Find the value of universal gravitational constant in cgs system

Grade:11

1 Answers

Arun
6 years ago
 

in MKS units G=6.67*10-11 Nm2/kg2, one N=kgm/s2 =1000g*100cm/s2 =105dyne,1m^2=10^4 cm2, 1kg^2=10^6g2,

 

G=6.67*10^(-11)*105*104/10^6 =6.67*10-8dyne*cm^2/g^2